# MAPS: A Multi-Agent Framework Based on Big Seven Personality and Socratic Guidance for Multimodal Scientific Problem Solving

### v1

@@ 2.3 Four-Step MSPs Solving
Table 1: Performance of different models across 10 subtasks from 3 datasets. ‘Gen.’ refers to General problems in the Mathvista dataset, ‘Phy.’ and ‘Chem.’ denote Physics and Chemistry tasks in the EMMA dataset. ‘MECO’, ‘MZCO’, and ‘MZCE’ correspond to COMP problems in English, COMP problems in Chinese, and CEE problems in Chinese, respectively, within the OlympiadBench dataset’s mathematics tasks. ‘PECO’ and ‘PZCE’ represent the COMP problems in English and CEE problems in Chinese under the physics tasks of the OlympiadBench dataset.
Scholar. After obtaining the semantically fused $l_{i}$, the model has essentially grasped the main idea of the problem. However, for scientific questions, mere comprehension is insufficient, as they typically require profound domain-specific knowledge, which is crucial for problem-solving. As illustrated in Figure 1, certain problems necessitate mastery of specialized formulas, such as those related to the principles of lever balance. Therefore, the Scholar agent is responsible for acquiring and integrating domain-specific knowledge based on $l_{i}$, thereby providing the necessary theoretical and practical support for the final solution $s_{i}$:
$s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i}),$ (4)
where $\mathcal{M}_{\text{sch}}(\cdot)$ represents the Scholar agent.
Solver. After the three steps involving the Interpreter, Aligner, and Scholar agents, the system has extracted and fused extensive information from both visual and textual modalities, providing rich contextual and specialized knowledge for the task. However, the real challenge lies in synthesizing this fragmented information into a coherent and logically sound solution. The fourth step, carried out by the Solver agent, bears the greatest pressure. This agent must not only integrate the information gathered in previous steps but also engage in deep reasoning to derive the final answer. With access to the most comprehensive data, the Solver has a unique advantage, allowing it to simplify complex issues and ensure the final answer $a_{i}$ is accurate and meets the expected standards, as given by:
$a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i}),$ (5)
where $\mathcal{M}_{\text{sol}}(\cdot)$ represents the Solver agent.
2.4 Critic and Feedback
Critic. This section introduces the Critic agent, which is responsible for evaluating and providing feedback on the four-step MSPs solving process. Rooted in Socratic questioning, the Critic prompts agents to reflect on their reasoning by asking questions such as “What assumptions are you making?” or “How can you justify this decision?” This approach encourages critical thinking and self-examination, key aspects of the Socratic method. Rather than relying on the final solution labels, the Critic evaluates the logic and justification of each step. The goal is to uncover flaws and push agents to refine their thought processes. Based on this feedback, the system identifies the weakest step and initiates a rollback and a redo to enhance the problem-solving process. The formula is given by:
$r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i}),$ (6)
where $\mathcal{M}_{\text{crit}}(\cdot)$ represents the Critic agent. $r_{i}$ is the scores and assessments of the four steps.
3 Experiments
We conduct extensive experiments on three MSPs datasets. This section introduces the three MSPs datasets and baselines in Section 3.1, and presents the experimental results and analysis in Section 3.2. The analysis and statistics of the Critic agent are presented in Section 3.3.
3.1 Datasets and Baselines
This study uses the latest three MSPs datasets: MathVista [23], OlympiadBench [16], and EMMA [14]. MathVista focuses on mathematics and general domains to assess models’ ability to solve complex scientific problems; OlympiadBench targets mathematics and physics to evaluate performance on Olympiad-level challenges; and EMMA, a multimodal dataset covering mathematics, physics, and chemistry, tests interdisciplinary scientific reasoning. Detailed descriptions of the tasks and datasets are in Appendix A, along with domain abbreviations, full names, and dataset descriptions in Table 1.
In the experimental setup, we select GPT-4o [^2] https://openai.com/api/ as our agent. For the baselines, we incorporate the current mainstream MLLMs, both closed-source and open-source, and evaluated them in both direct and CoT settings. The experimental settings and baselines are detailed in Appendix B.
3.2 Main Results
MAPS has set a new state-of-the-art (SOTA) performance, surpassing human-level performance for the first time. The results in Table 1 provide a detailed comparison of MAPS and the baselines. As shown, MAPS outperforms the current SOTA by 15.84% across all tasks, demonstrating its superiority in MSPs problem-solving. Additionally, MAPS slightly exceeds human experts by 3.58% in overall performance. MAPS excels in mathematical, physical, chemical, and general problems, demonstrating strong interdisciplinary reasoning. The multi-agent system, representing the Big Seven personalities, sets a new SOTA in collaboration. Experimental results show MAPS’s advantage in multimodal semantic fusion and multi-step reasoning, leveraging its ability to integrate diagrams with context and questions. This boosts problem-solving by enhancing domain-specific knowledge. The Critic agent, using Socratic feedback, offers suggestions for improvement, improving accuracy and reliability in complex tasks.
@@ 3.2 Main Results
In the MathVista dataset, the problem types are diverse, covering judgment questions, multiple-choice questions, and open-ended fill-in-the-blank questions, with various answer types such as text, floats, and integers. This requires the system to possess specialized capabilities, enabling it to accurately understand the question’s intent and provide specific responses. In well-known open-ended problems like those in OlympiadBench, the solution process involves complex steps and high difficulty, with open-ended answers that can lead to completely different outcomes due to subtle changes. In the EMMA dataset, the questions not only involve diagrams, but some options also contain diagrams, significantly increasing the challenge of understanding and aligning multiple diagrams. Furthermore, the question types are equally diverse, including open-ended and multiple-choice questions, raising the demands on the solving methods. The MAPS framework, through feedback-based multi-agent collaborative learning and the use of Socratic questioning, successfully addresses these different types of sub-tasks, achieving the latest SOTA. It demonstrates MAPS’s perfect adaptability and robustness when dealing with the diversity of question types, task difficulties, and multi-step reasoning.
3.3 Analysis of Critic Agent
The Critic agent plays a crucial feedback role in MAPS, enhancing the overall reasoning process without the reliance on gold labels. As shown in Figure 4, the upper part introduces the critic and feedback schema, while the lower part presents the feedback proportions provided by the Critic agent to each of the agents across three datasets, visually illustrating the feedback and corrections made by each assistant agent. The Critic agent’s feedback mechanism is rooted in Socratic questioning, which encourages deep reflection and rethinking of the reasoning steps. Instead of simply evaluating the process, the Critic fosters critical thinking, prompting the system to question its assumptions and improve the approach.
The schema of the Critic agent consists of two components: scores and feedback. The scores record the ratings given by the Critic agent for the four problem-solving steps, ranging from 0 to 5. Higher scores indicate better execution, while lower scores highlight areas needing improvement. Feedback primarily captures the Critic agent’s Socratic-style, heuristic suggestions designed to encourage deeper reflection on each step. These suggestions are not limited to evaluations; they also guide the agents to reconsider and correct their reasoning process. For example, the Critic may prompt the agents to think about alternative approaches or reassess their conclusions, challenging them to think more critically. To ensure the final result’s quality, a rule is set in the Manager agent: if any step receives a low score, the system will backtrack and make corrections. If all steps score well, the system outputs the result directly. This approach guarantees the effectiveness of each step, optimizing the overall problem-solving process.
In the EMMA and OlympiadBench datasets, the Solver agent exhibits the highest feedback proportion. The lower part of Figure 4 shows the distribution of agents that performed feedback corrections across the three datasets. It can be observed that the feedback proportions vary across the three datasets. In the MathVista dataset, the highest feedback proportion corresponds to no regeneration, indicating that the overall solution performance on this dataset is still quite good. This result is consistent with the findings in Table 1, and we have surpassed the previous SOTA by a 5.0% improvement. This suggests that while there is still room for further enhancement in certain areas, the system’s overall reasoning capability is already quite impressive when dealing with more complex problems. In contrast, in the EMMA and OlympiadBench datasets, the Solver agent shows the highest feedback proportion, suggesting that this part is the most difficult and prone to errors in these datasets. Specifically, this part includes the first three steps: interpretation, alignment, and integration of information provided by the Scholar. These steps require the integration of a large amount of data, involve the most complex reasoning, and bear the most critical responsibility—providing a complete and accurate solution to the problem. In comparison, the feedback proportions from the other three agents are similar, indicating that their error rates are relatively low, and the performance of these sub-steps is relatively better.
4 Supplementary Analysis
To conduct more comprehensive experiments, this section covers the following topics: ablation experiments in Section 4.1, the base model generalization in Section 4.2, the time efficiency of MSPs solving in Section 4.3.
@@ 4 Supplementary Analysis
Due to space limitations, generalization experiments on other datasets, such as DiagramQG, are discussed in Appendix C. The case study and complete process for solving the MSPs task is outlined in Appendix D, and the prompts for all agents are provided in Appendix E.
4.1 Ablation Study
@@
Ablating the Interpreter results in the greatest loss of performance. We conduct ablation experiments on the OlympiadBench dataset to evaluate the impact of each module on the overall performance.  Table 2 presents the effects of removing the Interpreter, Aligner, Scholar, and Critic modules from the MAPS framework. The results show that removing the Interpreter agent causes the largest performance degradation, at 16.09%. This is because, in MSPs tasks, diagrams contain a wealth of valuable information, which serves as an important supplement to the text. Understanding diagrams plays a crucial role in problem-solving.
@@
The removal of the Critic agent causes the smallest performance loss. It results in only a 7.05% decrease, underscoring its role in providing feedback and corrections. While this mechanism allows MAPS to backtrack and refine its reasoning, its impact is less significant than that of other agents. Removing the Scholar agent results in 11.49% performance drops, highlighting the importance of searching and integrating domain-specific knowledge. Finally, the removal of the Aligner agent causes a 10.86% drop, indicating that while diagram and context alignment is valuable, its effect is smaller compared to other components.
4.2 Base Model Generalization
MAPS can improve performance across different base models. We conduct a series of experiments to verify whether our MAPS framework can demonstrate robust generalization across different base LLMs. The results showcase the model’s robustness and transferability, further affirming the high adaptability of the MAPS framework, which achieves consistently excellent performance across various foundational models. In particular, this underscores the universality and flexibility of MAPS. To further validate its generalization across different models, we specifically evaluate another base model, Qwen2.5-VL-72B, as well as Gemini 2.0 Flash, thereby demonstrating that MAPS performs outstandingly across models of varying scales and capabilities.  Figure 5 presents the experimental results of three sets of base models. In each set, the performance of mathematical, physical, and chemical sub-tasks is compared between MLLMs and MAPS with the respective base model. The results show varying degrees of improvement for each task. For example, in the physical task, $MAPS_{Qwen}$, based on Qwen2.5-VL-72B, shows a 12.4% improvement compared to Qwen2.5-VL-72B, while $MAPS_{Gemini}$, based on Gemini, shows a 4.2% improvement compared to Gemini. Similar improvements are observed in the mathematical and chemical sub-tasks as well. This indicates that MAPS can extend effectively on both closed-source and open-source MLLMs and achieve good experimental results across different base models of MLLMs.
@@ 4.2 Base Model Generalization
4.3 Time Efficiency
Solving time efficiency varies by question type, answer type, category, and difficulty, with multiple-choice and integer-type questions being the fastest, while higher difficulties and more complex questions require more time.  Figure 6 presents the solving time efficiency under different question types, answer types, categories, and difficulties, with all times represented relative to a 100s benchmark.
From the graphs, it is clear that multiple-choice questions are solved faster as the options are already provided, making the process more time-efficient. In contrast, open-ended questions take longer due to the lack of predefined answers and the need for deeper reasoning and analysis. Integer-type answers show the highest efficiency, as they often correspond to simpler questions or multiple-choice tasks that involve less reasoning. General questions also have the shortest solving times, as they typically require less complex reasoning compared to more specialized problem types. Finally, higher difficult levels lead to a decrease in solving time efficiency, as more complex problems require more time for deeper reasoning and analysis.
5 Related Works
@@
The related work is structured into two main aspects: first, an introduction to MSPs solving; and second, an exploration of multi-agent techniques.
@@
MSPs Solving. The research of MSPs spans across multiple fields, including mathematics, physics, and chemistry, with each area focusing on enhancing problem-solving abilities. In mathematics, studies [9, 11, 34] explore various methods to improve mathematical problem-solving, such as algorithm optimization, educational strategies, and the use of artificial intelligence. These approaches aim to boost the efficiency, accuracy, and depth of mathematical reasoning, fostering innovation and progress in the field. In the field of physics, the papers [24, 17, 3] emphasize the integration of different information types, such as images and text, through multimodal learning to enhance the efficiency and precision of problem solving. By utilizing techniques like reinforcement learning, generative models, and intelligent prompting, these studies optimize models’ performance in tackling complex physics problems. In chemistry, three articles [1, 18, 21] investigate the role of multimodal learning in solving chemical problems. By combining diverse information sources, including images and text, and employing techniques such as generative models and molecular geometry reasoning, they aim to improve both the efficiency and accuracy of solving chemistry problems, driving innovation in chemistry education and research.
@@
Multi-Agent. Multi-agent systems, built on LLMs, consist of multiple AI agents that specialize in specific tasks, working together to solve complex problems [30, 42, 39, 32]. When presented with a problem, these agents decompose it into smaller, manageable subtasks and utilize various tools, such as internet data retrieval, to solve them through iterative steps. Several studies [28, 37, 40, 26] have employed multi-agent systems to tackle challenges like problem identification, code writing and debugging, data visualization, and providing interactive feedback to human users. In their work, Ni and Buehler [25] highlights the potential of AI-driven multi-agent teams in solving mechanical problems autonomously, demonstrating an enhanced capability for understanding, formulating, and validating engineering solutions through self-correction and collaborative refinement. Inspired by the research, we developed the MAPS method, which leverages multi-agent collaborative learning and stepwise problem-solving to provide innovative solutions for MSPs. By combining the strengths of AI agents, complex problems can be broken down into subtasks and solved step by step through collaboration, improving efficiency and accuracy.
6 Conclusion
@@
This study introduces a MAPS approach utilizing a multi-agent framework based on the Big Seven Personality theory and Socratic guidance to tackle the challenges of multi-modal comprehensive reasoning and the lack of reflective capabilities. The framework involves seven agents, each specializing in distinct aspects of problem-solving. To address the first challenge, a four-agent strategy is proposed, where each agent focuses on specific stages of the reasoning process. Additionally, the Critic agent addresses the second challenge through Socratic reflection and critical feedback. Extensive experiments on the EMMA, Olympiad, and MathVista datasets validate MAPS’s effectiveness in overcoming these issues and enhancing performance across various reasoning tasks. Meanwhile, we perform additional analytical experiments to assess the model’s advancement as well as its generalization.
Supplementary Material
Appendix A Datasets and Tasks

### v2

@@ Introduction
Table 1: Performance across 10 subtasks. Gen. = General (MathVista), Phy./Chem. = Physics/Chemistry (EMMA), MECO/ MZCO/ MZCE = English/Chinese COMP & CEE Math (OlympiadBench), PECO/ PZCE = English/Chinese Physics (OlympiadBench).
Critic and Feedback
The Critic agent evaluates the internal reasoning trajectory $\mathcal{T}=\{p_{i},l_{i},s_{i},a_{i}\}$ without relying on ground-truth answers. Inspired by Socratic questioning, it examines each stage’s logic and justification to identify flawed assumptions and incomplete inferences.
We define the feedback vector as:
$\mathbf{f}_{i}=\mathcal{M}_{\text{crit}}(p_{i},l_{i},s_{i},a_{i}),\quad\mathbf{f}_{i}\in[0,1]^{4},$ (12)
where each element $f_{i}^{(k)}$ represents the Critic’s confidence in the correctness and completeness of stage $k$.
The weakest stage is selected by:
$k^{\ast}=\arg\min_{k}f_{i}^{(k)},\quad\text{if}\quad f_{i}^{(k^{\ast})}<\tau,$ (13)
which triggers a targeted revision. This reflection-driven loop promotes iterative self-correction and deepens reasoning reliability.
Experiments
Datasets and Baselines
We evaluate on three benchmarks for complex problem reasoning: MathVista, OlympiadBench, and EMMA. MathVista covers math and general science, OlympiadBench focuses on high-level math and physics, and EMMA includes math, physics, and chemistry. See Appendix B for details.
We use GPT-4o (achiam2023gpt) as the base model. For comparison, we include leading multimodal large language models (MLLMs), both proprietary and open-source, tested under both direct and Chain-of-Thought (CoT) settings. Experimental details and baseline setups are provided in Appendix C.
Main Results
MAPS achieves a new state-of-the-art (SOTA), surpassing human-level performance for the first time.
As shown in Table 1, MAPS outperforms previous SOTA models by 15.84% and exceeds human expert performance by 3.58% across all tasks, highlighting its strength in solving complex multimodal problems. MAPS demonstrates robust performance across mathematical, physical, chemical, and general tasks, showcasing strong interdisciplinary reasoning. Its multi-agent design, based on the Big Five personality theory, enables effective collaboration and contributes to the SOTA results. The system excels in multimodal semantic integration and multi-step reasoning by jointly leveraging diagrams, contexts, and questions. Furthermore, the Critic agent applies Socratic feedback to refine responses, enhancing both accuracy and reliability on challenging tasks.
MAPS exhibits strong adaptability and robustness across diverse reasoning tasks.
The evaluation datasets span a wide range of question types, modalities, and reasoning difficulties. MathVista features judgment, multiple-choice, and open-ended fill-in-the-blank questions with varied answer formats, requiring accurate intent understanding and response generation. OlympiadBench emphasizes challenging open-ended problems demanding multi-step symbolic reasoning, where small errors can lead to divergent outcomes. EMMA introduces multimodal complexity with diagrams embedded in both questions and answer choices. Through feedback-driven multi-agent collaboration and Socratic questioning, MAPS effectively handles these challenges, achieving SOTA and demonstrating strong generalization across heterogeneous reasoning scenarios.
Analysis of Critic Agent
Critic schema includes scoring and Socratic feedback.
Each reasoning step is rated from 0–5, guiding whether the system should backtrack or proceed. Feedback is heuristic, encouraging rethinking rather than offering direct answers. The Critic uses these scores to decide whether to regenerate specific steps, ensuring robustness in the final answer.
The Critic enhances reasoning via Socratic feedback without using gold labels.
As shown in Figure 5, the top illustrates the feedback schema, while the bottom shows feedback proportions across three datasets. The Critic prompts reflection and correction by encouraging agents to question assumptions rather than passively accept reasoning.
Solver receives the most feedback in EMMA and OlympiadBench.
As shown in the lower half of Figure 5, feedback varies by dataset. For MathVista, most steps need no regeneration, aligning with our superior 5.0% SOTA improvement in Table 1. This reflects strong baseline reasoning. In contrast, EMMA and OlympiadBench show the highest feedback for the Solver, especially in the interpretation, alignment, and integration steps. These are the most complex and error-prone stages. Other agents receive comparable and lower feedback, indicating better relative performance in their sub-tasks.
@@
Supplementary Analysis
@@
Due to space limitations, generalization experiments on other datasets, such as DiagramQG (zhang2025diagram), are discussed in Appendix D. The case study and complete process of MAPS is outlined in Appendix E, and the prompts for agents are provided in Appendix F.
Ablation Studies
Ablating the Interpreter results in the greatest loss of performance.
We conduct ablation experiments on the OlympiadBench dataset to evaluate the impact of each module on the overall performance. Table 2 presents the effects of removing the Interpreter, Aligner, Scholar, and Critic modules from the MAPS framework. The results show that removing the Interpreter agent causes the largest performance degradation, at 16.09%. This is because, in complex problem reasoning, diagrams contain a wealth of valuable information, which serves as an important supplement to the text. Understanding diagrams plays a crucial role in problem-solving.
The removal of the Critic agent causes the smallest performance loss.
It results in only a 7.05% decrease, underscoring its role in providing feedback and corrections. While this mechanism allows MAPS to backtrack and refine its reasoning, its impact is less than that of other agents. Removing the Scholar agent results in 11.49% performance drops, highlighting the importance of searching and integrating domain-specific knowledge. Finally, the removal of the Aligner agent causes a 10.86% drop, indicating that while diagram and context alignment is valuable, its effect is smaller compared to other components.
@@
Base Model Generalization
MAPS improves performance across diverse base models.
We conduct experiments to verify whether our MAPS framework demonstrates robust generalization across various base LLMs. The results confirm MAPS’s robustness and transferability, highlighting its adaptability and consistent performance across different foundation models. To further validate its generalization, we evaluate both Qwen2.5-VL-72B and Gemini 2.0 Flash, showing that MAPS performs well across models of varying scales and capabilities. Figure 6 presents results for three sets of base models. In each set, we compare MLLMs and MAPS on mathematical, physical, and chemical sub-tasks. MAPS consistently outperforms the base models. For example, MAPSQwen improves Qwen2.5-VL-72B by 12.4% in physics, while MAPSGemini improves Gemini by 4.2%. Similar gains are observed in math and chemistry, demonstrating MAPS’s effectiveness on both open-source and closed-source MLLMs.
Time Efficiency
Simpler formats and lower difficulty yield faster solving times.
Solving time efficiency varies by question type, answer type, category, and difficulty, with multiple-choice and integer-type questions being the fastest, while higher difficulties and complex formats require more time. Figure 7 illustrates the solving time efficiency across various dimensions—question types, answer formats, subject categories, and difficulty levels—with all times normalized to a 100s benchmark.
Predefined structure and conceptual simplicity reduce reasoning time.
Multiple-choice questions are solved more quickly thanks to predefined answer options that limit the need for extensive reasoning or exploration. Integer-type answers also show high efficiency, often tied to simpler arithmetic or structured formats requiring minimal inference. General category questions are faster on average, likely due to lower conceptual and reasoning complexity compared to domain-specific tasks. In contrast, open-ended questions demand deeper analysis and justification, leading to longer solving times. Finally, solving efficiency declines with increased difficulty: as question complexity rises, so does the required reasoning time, reflecting greater cognitive and computational demands.
@@
Related Works
@@
The related work is structured into two main aspects: Firstly, an introduction to complex problem reasoning; Secondly, an exploration of multi-agent techniques.
Complex Problem Reasoning.
@@
The research of complex problem reasoning spans across multiple fields, including mathematics, physics, and chemistry, with each area focusing on enhancing problem-solving abilities. In mathematics, studies (didolkar2024metacognitive; fitriana2024influence; tong2025dart) explore various methods to improve mathematical problem-solving, such as algorithm optimization, educational strategies, and the use of artificial intelligence. These approaches aim to boost the efficiency, accuracy, and depth of mathematical reasoning. In the field of physics, the papers (mustofa2024utilizing; kapuriya2024mm; anand2024mm; wu2024gps) emphasize the integration of different information types, such as images and text, through multimodal learning to enhance the efficiency and precision of problem solving. In chemistry, three articles (alasadi2024multimodal; kiernan2024resources; li2024chemvlm; lin2025foundation; dang2025disentangled) investigate the role of multimodal learning in solving chemical problems. By combining diverse information sources, including images and text, and employing techniques such as generative models and molecular geometry reasoning, they aim to improve both the efficiency and accuracy of solving chemistry problems.
Multi-Agent.
@@
Multi-agent systems, built on LLMs, consist of multiple AI agents that specialize in specific tasks, working together to solve complex problems (richards2023auto; yang2023auto; wu2023autogen; sun2023corex; zhang2025maps; zhang2025mars; zhang2025gkg; yan2025mur). When presented with a problem, these agents decompose it into smaller, manageable subtasks and utilize various tools, such as internet data retrieval, to solve them through iterative steps. Several studies (poldrack2023ai; wang2024survey; xi2025rise; ni2021deep) have employed multi-agent systems to tackle challenges like problem identification, code writing and debugging, data visualization, and providing interactive feedback to human users. In their work, ni2024mechagents highlights the potential of AI-driven multi-agent teams in solving mechanical problems autonomously, demonstrating an enhanced capability for understanding, formulating, and validating engineering solutions through self-correction and collaborative refinement. Inspired by the research, we developed the MAPS method, which leverages multi-agent collaborative learning and stepwise problem-solving to provide innovative solutions for complex problem reasoning. By combining the strengths of AI agents, complex problems can be broken down into subtasks and solved step by step through collaboration, improving efficiency and accuracy.
Conclusion
@@
This study presents MAPS, a multi-agent framework grounded in the Big Five Personality Theory and guided by Socratic principles, designed to address the challenges of multimodal comprehensive reasoning and enhance reflective capabilities. The framework involves five agents, each specializing in distinct aspects of problem-solving. To address the first challenge, a four-agent strategy is proposed, where each agent focuses on specific stages of the reasoning process. Additionally, the Critic agent addresses the second challenge through Socratic reflection and critical feedback. Extensive experiments on the EMMA, OlympiadBench, and MathVista datasets validate MAPS’s effectiveness in overcoming these issues and enhancing performance across various reasoning tasks. Meanwhile, we perform additional analytical experiments to assess the model’s advancement as well as its generalization.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62137002, 62277042, 62293553, 62450005, 62437002, 62477036, 62477037, 62176209, 62192781, 62306229), the “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), and the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”.
Appendix A Proof of Proposition
Proof of Proposition 1 (Monotonic Free-Energy Descent)
Statement recap. Let $F^{(t)}$ denote the variational free energy after the $t$-th MAPS iteration:
$F^{(t)}=\mathbb{E}_{q^{(t)}(\theta)}[-\log p(\mathbf{x},a_{i}\mid\theta)]+\mathrm{KL}(q^{(t)}(\theta)\|p(\theta)),$ (14)
where $q^{(t)}(\theta)$ is the approximate posterior induced by the agent outputs at iteration $t$. The proposition claims that each Critic-guided update ensures
$F^{(t+1)}\leq F^{(t)},$ (15)
which implies that $\{F^{(t)}\}_{t=0}^{\infty}$ is a non-increasing sequence that converges.
Step 1: Free Energy as a Variational Upper Bound. By standard variational inference, the marginal log-likelihood satisfies
$\log p(\mathbf{x},a_{i})=\mathcal{L}(q)+\mathrm{KL}(q(\theta)\|p(\theta\mid\mathbf{x},a_{i})),$ (16)
where the ELBO is defined as
$\mathcal{L}(q)=\mathbb{E}_{q(\theta)}[\log p(\mathbf{x},a_{i}\mid\theta)]-\mathrm{KL}(q(\theta)\|p(\theta)).$ (17)
This leads to the variational free energy $F(q):=-\mathcal{L}(q)$, which upper-bounds the negative log marginal likelihood:
$F(q)\geq-\log p(\mathbf{x},a_{i}).$ (18)
Step 2: Critic-Guided Local Update. Suppose the Critic identifies the weakest stage $k^{*}=\arg\min_{k}f_{k}^{(t)}$ and re-executes $\mathcal{A}_{k^{*}}$, yielding updated output $\mathcal{S}_{k^{*}}^{(t+1)}$. The updated trajectory leads to a refined posterior:
$q^{(t+1)}(\theta)=\mathcal{G}(q^{(t)}(\theta),\mathcal{S}_{k^{*}}^{(t+1)}),$ (19)
where $\mathcal{G}$ represents the update mechanism that propagates changes forward through dependent modules.
As the re-computed stage corrects suboptimal reasoning or improves knowledge alignment, it leads to a strictly better or equal ELBO, and thus
$F^{(t+1)}=F(q^{(t+1)})\leq F(q^{(t)})=F^{(t)}.$ (20)
Step 3: Convergence. Since $F^{(t)}$ is lower-bounded and non-increasing, the sequence converges:
$\lim_{t\to\infty}F^{(t)}=F^{(\infty)}.$ (21)
Proof of Proposition 3 (Collaborative Information Bottleneck)
Statement recap. Given input $\mathbf{x}=(d_{i},c_{i},q_{i})$ and target answer $a_{i}$, let $\mathcal{S}_{k}$ denote the intermediate output of the $k$-th agent in the MAPS reasoning process. Then the objective is to find agents that minimize cumulative information compression from input while preserving task-relevant content:
$\min_{\{\mathcal{A}_{k}\}}\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad I(\mathcal{S}_{k};a_{i})\geq\varepsilon,$ (22)
where $I(\cdot;\cdot)$ denotes mutual information, and $\varepsilon>0$ is a threshold for information sufficiency.
Step 1: Motivation from the Information Bottleneck (IB) Principle. In standard IB theory (tishby2000information), one seeks a compressed representation $\mathcal{S}$ of input $\mathbf{x}$ that retains information about target $a_{i}$. The trade-off objective is:
$\min I(\mathbf{x};\mathcal{S})\quad\text{subject to}\quad I(\mathcal{S};a_{i})\geq\varepsilon.$ (23)
MAPS generalizes this to a multi-agent, multi-stage setting, where each agent $\mathcal{A}_{k}$ produces $\mathcal{S}_{k}$, and the full reasoning trajectory is $\mathcal{T}=\{\mathcal{S}_{1},\dots,\mathcal{S}_{4}\}$. The objective thus becomes collaborative:
$\min\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad\forall k,\;I(\mathcal{S}_{k};a_{i})\geq\varepsilon.$ (24)
This reflects two goals: (1) compress redundant input features per stage, and (2) preserve task-relevant evidence at each step.
Step 2: Role of the Critic as a Constraint Monitor. The Critic observes the reasoning trajectory $\mathcal{T}$ and computes scores $\mathbf{f}=\{f_{1},f_{2},f_{3},f_{4}\}$, where each $f_{k}$ implicitly estimates $I(\mathcal{S}_{k};a_{i})$. If $f_{k}<\tau$, we assume the sufficiency constraint is violated:
[…]
The revision process updates $\mathcal{S}_{k}$ to $\mathcal{S}_{k}^{\prime}$, ideally increasing $I(\mathcal{S}_{k}^{\prime};a_{i})$ while not increasing $I(\mathbf{x};\mathcal{S}_{k}^{\prime})$. Since the Critic only intervenes when constraints are violated, each update maintains feasibility and possibly reduces the objective.
Step 3: Convergence and Validity. As each update either preserves or improves the sufficiency $I(\mathcal{S}_{k};a_{i})$, and the overall input compression is bounded below by zero, the process converges to a locally optimal set of agents $\{\mathcal{A}_{k}^{*}\}$ satisfying:
$\forall k,\quad I(\mathcal{S}_{k}^{*};a_{i})\geq\varepsilon,\quad\text{with minimal}\quad\sum_{k}I(\mathbf{x};\mathcal{S}_{k}^{*}).$ (26)
This completes the proof that the MAPS four-step process can be viewed as a constrained collaborative information bottleneck optimization.
Appendix B Datasets